OpenStudy (anonymous):

Boating on a lake. The Clarke family went sailing on a lake. Their boat averaged 6 kilometers per hour. The Rourke family took their outboard run about for a trip on the lake for the same amount of time. Their boat averaged 14 kilometers per hour. The Rourke family traveled 20 kilometers father than the Clarke family. How many hours did each family spend on their boat trip?

OpenStudy (anonymous):

2.5 Hours, since (14h-6h)=20 so, 14*2.5-6*2.5=20 -> 35-12 = 20 km.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem kiddo.